OpenStudy (anonymous):

what is the theoretical probability of rolling a sum of 8 on one roll of two standard number cubes?

OpenStudy (anonymous):

26,35,44,53,62=?

OpenStudy (loser66):

8 = 2+6, right? 8 = 3 +5, right? 8 = 4+4, right? when drawing out the grids, how many grids are there on those forms?

OpenStudy (loser66):

 there are some more, you find them all, count how many grids they occupied then, divide by 36